SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Sept 13, 2005)

Attendees: Kris Chaisanguanthum, Al Eisner, Fang Fang, Larry Gladney, Selina Li, Dong Su

Operations

  • The old DCT system is taking out of data-taking starting from today.
  • Two L1Passthru physics runs were taken last Saturday. See HN posted.

EMT

  • Rainer helped Dan debugging the EMT prescale problems encountered when testing reading the configuration xtc from the ROM lib. It turned out that Dan needed to link codes related to the the prescale in the makefile. However, there is still one more error remains to be solved.

DCT/DCZ

  • Last Wednesday, Selina tested powering off the new DCZ crates and see what happened. For more detailed info, see HN posted. Jim Hamilton pointed out that the new DCZ crates can be put in the daqCrateMonitor Proxy so that the run control will stop taking data if there is something wrong with one of these trigger crates. This is in Jim's to-do-list.
  • Today, Rainer updated the crate mask to exclude the old DCT from data-taking.
  • Su Dong pointed out that MPIC spurious interruptions were seen in the TSF-X ROM logfile during run 57958 (see HN posted).

GLT

  • Su Dong looked at the trigger rate for the 1E run taken last Saturday, and found out that about 180 Hz is exclusively for 1E, that is, 1E trigger contributes about 10% of the total trigger rate. According to Su Dong, L1 needs additional logic to make us comfortable in triggering for single photon physics, and L3 need special filter to log as much data as they can. Su Dong is looking at data at the trigger side to see if there are any hints to help reduce the rate by at least a factor of 2. The single phone physics group also needs to demonstrate that the physics is feasible.

L3

  • Nothing and no one to report.

Data Quality and Monitoring

  • Kris took care of the old DCT related plots for fastmon and JAS. New version of fastmon is running for production.
  • JAS couldn't disply HepMultiScaler properly (see examples here for Ap,B and event rate ).
  • Run 57965 has incorrect time range for the time history plots, and after this run, the .ps archive was stopped updating. Kris will ask Jim what happened.

Online

  • Fang talked about the dependency of the DCT codes for IR2 and teststand. See the slides here.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.